# Magnetic tunnel junction equivalent circuit modeling

### v2

@@ 3 Validation of Equivalent Circuit Model
The circuit shown in Fig. 1(b) was implemented in LTspice, a widely used free simulation tool available at [Analog Devices]. This model simulates MTJ dynamics under time-dependent voltages $V_{\mathrm{MTJ}}$ and magnetic fields $\mathbf{B}_{e}$. This section presents simulations of ferromagnetic resonance (FMR), operation of the MTJ as a spin-torque nano-oscillator (STNO), field-induced switching, and spin-torque switching. The LTspice model and all simulation files are available at [MTJ Spice Models].
@@ 3.1 Ferromagnetic Resonance
The FMR response was evaluated by comparing LTspice simulations of the equivalent circuit model with theoretical predictions. The simulations modeled free decay of magnetization oscillations, i.e., the magnetization was initialized slightly off equilibrium and then relaxed with oscillations at the FMR frequency. Simulations used a free layer with magnetization $M_{s}=796$ kA/m ($B_{d}=\mu_{0}M_{s}=1$ T), volume of $V_{s}=5.65\times 10^{-24}$ m3, corresponding to an elliptical geometry of the free layer with a semi-axes of 30 nm and 20 nm and a thickness of 3 nm, and the anisotropy field $B_{a}=0.2$ T. To precisely measure the FMR frequency, the Gilbert damping parameter was set to a small value of $\alpha_{G}=0.0001$.
@@
The agreement between simulated FMR frequencies and theoretical values confirms the accuracy of the equivalent circuit. The extremely low error values demonstrate the validity of the circuit model for external magnetic fields oriented along the $\hat{\mathbf{x}}$, $\hat{\mathbf{y}}$, and $\hat{\mathbf{z}}$ axes, including cases involving anisotropy and mode transitions (e.g., when $B_{e,y}=B_{a}$).
@@ 3.2 Spin-Torque Nano-Oscillator
STNOs are nanoscale devices that use spin-transfer torque to generate self-sustained oscillations in the free layer magnetization, which produce a time-varying resistance in the MTJ [Slavin 2009]. Here, we demonstrate that the equivalent circuit model accurately reproduces large-amplitude oscillations in a durrent-driven MTJ.
Figure 3(a) shows the trajectory of the free layer magnetization, $\mathbf{m}$, simulated in LTspice with a spin-polarized current of 0.33 mA, $R_{\mathrm{P}}=500~{}\Omega$, $R_{\mathrm{AP}}=1500~{}\Omega$, spin polarization vector oriented at $\theta_{p}=85^{\circ}$ and $\phi_{p}=190^{\circ}$, Gilbert damping constant $\alpha_{G}=0.01$, anisotropy field of $B_{a}=0.2$ T, and external magnetic field components $B_{e,x}=-0.1$ T, $B_{e,y}=-0.15$ T, and $B_{e,z}=0.8$ T. The free layer magnetization undergoes sustained precessional motion at approximately 14 GHz, characteristic of STNO operation. The trajectory has non-elliptical shape due to the combined effects of the effective field and spin-transfer torque.
@@
Figures 3(b) presents the time evolution of the polar angle, $\theta$ and demonstrates that LTspice model produces results identical to LLGS equation. This close agreement demonstrates that the equivalent circuit accurately models all three components of the LLGS equation: the conservative precession term, the damping term, and the spin-transfer anti-damping torque. The simulations specifically employed unconventional external magnetic fields and polarization angles designed to produce non-uniform and non-elliptical precession; despite this complexity, the equivalent circuit performed well. These results may find application in integrated STNO-based systems [Litvinenko 2022], where accurate circuit-level simulation of complex magnetization dynamics would be a valuable tool for designing complete circuit implementations.
@@ 3.3 Field-Induced Switching
In Fig. 4(a) we compare LTspice and LLGS simulations for field-induced switching of an MTJ. In this simulation, the system was subjected to a time-varying external magnetic field applied for 0.058 ns at intervals of 1.0 ns, with the field turned off otherwise. When active, the field had an amplitude of 0.4 T and was oriented in the $xy$ plane at an angle $15^{\circ}$ (green shading in Fig. 4(a)) or $195^{\circ}$ (blue shading) to MTJ easy axis $\hat{\mathbf{x}}$. Other simulation parameters are $R_{\mathrm{P}}=500~{}\Omega$, $R_{\mathrm{AP}}=1500~{}\Omega$, the anisotropy field of $B_{a}=0.1$ T, Gilbert damping coefficient of $\alpha_{G}=0.05$.
The strong agreement between LTspice and LLGS results, including accurate description of post-switching “ringing”, confirms the accuracy of the equivalent circuit model for large-angle ($180^{\circ}$) magnetization rotation.
The previous two subsections demonstrated that the equivalent circuit functions accurately in stationary regimes, where magnetic fields and currents remain static while the output dynamically evolves. This section has shown that the circuit also performs reliably under time-varying magnetic fields. The example of field-induced switching was selected for its familiarity to the MRAM community, but additional tests (not shown) included a variety of time-dependent fields, such as sinusoidal and nonlinear impulses. In every case, the circuit produced results that closely matched macrospin simulations. Moreover, these simulations were straightforward to configure and fast to execute using the equivalent circuit in LTspice.
@@ 3.4 Current-Induced Switching in a 1T-1MTJ Cell
To further validate the equivalent circuit model and demonstrate the utility of LTspice simulations, the MTJ was integrated as a single LTspice component, available at [MTJ Spice Models 2025]. This integrated MTJ is highlighted by a yellow box in Figure 4(b), which shows a schematic of a 1T-1MTJ memory cell implemented in LTspice. A 1T-1MTJ cell consists of a single transistor and an MTJ connected in series, forming a compact unit suitable for high-density memory arrays [Guo 2015]. In this configuration, the wordline activates the write operation, while the bitline and source line provide the necessary bias and read voltages for proper cell operation. Current-induced switching occurs when an electric current reverses the magnetization of the free layer, toggling the MTJ between its low- and high-resistance states.
To evaluate the accuracy of the equivalent circuit model, current-induced switching was simulated in the 1T-1MTJ cell and compared to direct numerical simulation of the LLGS equation (see Fig. 4(c)). The excellent agreement between the two approaches confirms that the circuit model accurately reproduces the dynamics of current-induced switching.
In these simulations, a 5 V square wave with a period of 2 ns was applied alternately to the bitline and source line; the sourceline was at 5 V from 0 to 1 ns, and the bitline was at 5 V from 1 to 2 ns. Additionally, a 5 V square impulse lasting 0.1 ns was applied to the wordline at times 0.5 ns, 1.5 ns, and 2.5 ns. These applied voltage waveforms resulted in MTJ currents of approximately 10 mA at 0.5 ns and 2.5 ns (green shading in Fig. 4(c)) and –10 mA at 1.5 ns (blue shading). The spin polarization vector $\mathbf{p}$ was oriented in the $xy$ plane at $15^{\circ}$ relative to the $\hat{\mathbf{x}}$ axis. Other simulation parameters were the same as in the field switching example.
This section demonstrates the potential of the equivalent circuit model for simulating hybrid circuits with MTJs and conventional electronic components. Although a relatively simple 1T-1MTJ cell was chosen for the example, the equivalent circuit is applicable to circuits of any complexity. The equivalent circuit has been integrated into a single LTspice part, and readers are invited to download and incorporate this part into their own simulations.
@@ 4 Conclusion
In this work, we presented a physics-based equivalent circuit model for MTJs that accurately describes both magnetization dynamics and electrical characteristics. The proposed model was validated against direct numerical solutions of the LLGS equation across multiple operating regimes, including FMR, STNO, field- and current-induced switching. LTspice simulations demonstrated excellent agreement with numerical results, confirming the robustness and fidelity of the model.
@@
This work bridges spintronic physics and circuit-level implementation, providing an effective and efficient tool for the analysis and design of spintronic devices. The results presented here support the development of next-generation non-volatile memory and advanced computing technologies based on spintronic elements.

### v3

@@ 3 Validation of Equivalent Circuit Model
To assess the accuracy and practical utility of the proposed model, we performed a series of simulations across representative MTJ operating regimes. The simulated behaviors include ferromagnetic resonance (FMR), spin-torque nano-oscillator (STNO) operation, field-induced switching, and spin-torque switching.
@@ 3.1 Ferromagnetic Resonance
The FMR response was evaluated by comparing LTspice simulations of the equivalent circuit model with theoretical predictions. The simulations modeled free decay of magnetization oscillations, i.e., the magnetization was initialized slightly off equilibrium and then relaxed with oscillations at the FMR frequency. Simulations used a free layer with magnetization $M_{s}=796$ kA/m ($B_{d}=\mu_{0}M_{s}=1$ T), volume of $V_{o}=5.65\times 10^{-24}$ m3, corresponding to an elliptical geometry of the free layer with a semi-axes of 30 nm and 20 nm and a thickness of 3 nm, and the anisotropy field $B_{a}=0.2$ T. To precisely measure the FMR frequency, the Gilbert damping parameter was set to a small value of $\alpha_{G}=0.0001$.
@@
The agreement between simulated FMR frequencies and theoretical values confirms the accuracy of the equivalent circuit. The extremely low error values demonstrate the validity of the circuit model for external magnetic fields oriented along the $\hat{\mathbf{x}}$, $\hat{\mathbf{y}}$, and $\hat{\mathbf{z}}$ axes, including cases involving anisotropy and mode transitions (e.g., when $B_{e,y}=B_{a}$). These results indicate that the equivalent circuit accurately represents both the conservative (precessional) and damping components of the LLGS equation.
@@ 3.2 Spin-Torque Nano-Oscillator
STNOs are nanoscale devices that use spin-transfer torque to generate self-sustained oscillations in the free layer magnetization, which produce a time-varying resistance in the MTJ [Slavin 2009]. Here, we demonstrate that the equivalent circuit model accurately reproduces large-amplitude oscillations in a current-driven MTJ.
Figure 3(a) shows the trajectory of the free layer magnetization, $\mathbf{m}$, simulated in LTspice with $V_{\mathrm{mtj}}=0.2475$ V, $R_{\mathrm{P}}=500~{}\Omega$, $R_{\mathrm{AP}}=1500~{}\Omega$, spin polarization vector oriented at $\theta_{p}=85^{\circ}$ and $\phi_{p}=190^{\circ}$, Gilbert damping constant $\alpha_{G}=0.01$, anisotropy field of $B_{a}=0.2$ T, and external magnetic field components $B_{e,x}=-0.1$ T, $B_{e,y}=-0.15$ T, and $B_{e,z}=0.8$ T. The free layer magnetization undergoes sustained precessional motion at approximately 14 GHz. The trajectory has non-uniform shape due to the combined effects of the effective field orientation and STT orientation and $\theta_{p}=85^{\circ}$.
@@
Figures 3(b) presents the time evolution of the polar angle, $\theta$ and demonstrates that LTspice model produces results identical to LLGS equation. This close agreement demonstrates that the equivalent circuit accurately models all three components of the LLGS equation: the conservative precession term, the damping term, and the spin-transfer anti-damping torque. The simulations specifically employed unconventional external magnetic fields and polarization angles designed to produce non-uniform precession; despite this complexity, the equivalent circuit performed well. These results may find application in integrated STNO-based systems [Litvinenko 2022], where accurate circuit-level simulation of complex magnetization dynamics would be a valuable tool for designing complete circuit implementations.
@@ 3.3 Field-Induced Switching
In Fig. 4(a) we compare LTspice and LLGS simulations for field-induced switching of an MTJ. In this simulation, the system was subjected to a time-varying external magnetic field applied for 0.058 ns at intervals of 1.0 ns, with the field turned off otherwise. When active, the field had an amplitude of 0.4 T and was oriented in the $xy$ plane at an angle $15^{\circ}$ (green shading in Fig. 4(a)) or $195^{\circ}$ (blue shading) to MTJ easy axis $\hat{\mathbf{x}}$. Other simulation parameters are $R_{\mathrm{P}}=500~{}\Omega$, $R_{\mathrm{AP}}=1500~{}\Omega$, $B_{a}=0.1$ T, and $\alpha_{G}=0.05$.
The strong agreement between LTspice and LLGS results, including accurate depiction of post-switching “ringing”, confirms the accuracy of the equivalent circuit model for large-angle ($180^{\circ}$) magnetization rotation.
The previous two subsections demonstrated that the equivalent circuit functions accurately in stationary regimes, where magnetic fields and currents remain static while the output evolves dynamically. This section has shown that the circuit also performs reliably under time-varying magnetic fields. The example of field-induced switching was selected for its familiarity to the MRAM community, but additional tests (not shown) included various time-dependent fields, such as sinusoidal and nonlinear impulses. In every case, the circuit produced results that closely matched macrospin simulations. Moreover, these simulations were straightforward to configure and fast to execute using the equivalent circuit in LTspice.
@@ 3.4 Current-Induced Switching in a 1T-1MTJ Cell
To illustrate the practical application of the equivalent circuit model, the MTJ was incorporated into a 1T-1MTJ memory cell and simulated using LTspice. The schematic of this memory cell is shown in Fig. 4(b), and is as previously described in Fig. 1(d). A 1T-1MTJ cell comprises a single transistor in series with an MTJ, forming a compact and scalable memory element suitable for MRAM architectures [Guo 2015]. In this design, the wordline enables writing, while the bitline and source line supply the required voltages for read and write operations. Current-induced switching occurs when an electric current reverses the magnetization of the free layer, toggling the MTJ between low- and high-resistance states.
To evaluate the circuit model accuracy, current-induced switching was simulated in a 1T-1MTJ cell and compared with direct LLGS simulation (see Fig. 4(c)). The close agreement confirms the model faithfully reproduces current-induced switching dynamics.
In these simulations, a 5 V square wave with a period of 2 ns was applied alternately to the bitline and source line; the sourceline was at 5 V from 0 to 1 ns, and the bitline was at 5 V from 1 to 2 ns. Additionally, a 5 V square impulse lasting 0.1 ns was applied to the wordline at times 0.5 ns, 1.5 ns, and 2.5 ns. These applied voltage waveforms resulted in MTJ currents of approximately 9 mA at 0.5 ns and 2.5 ns (green shading) in Fig. 4(c)) and about –7 mA at 1.5 ns (blue shading). The spin polarization vector $\mathbf{p}$ was oriented in the $xy$ plane at $15^{\circ}$ relative to the $\hat{\mathbf{x}}$ axis. Other simulation parameters were the same as in the field switching example.
This section demonstrates the potential of the equivalent circuit model for simulating hybrid circuits with MTJs and conventional electronic components. Although a relatively simple 1T-1MTJ cell was chosen for the example, the equivalent circuit is applicable to circuits of any complexity. The equivalent circuit has been integrated into a single LTspice part, and readers are invited to download and incorporate this part into their own simulations [MTJ Spice Models].
@@ 4 Conclusion
In this work, we presented a physics-based equivalent circuit model for MTJs that accurately describes both magnetization dynamics and electrical characteristics. The proposed circuit model was validated against direct numerical solutions of the LLGS equation across multiple operating regimes, including FMR, STNO, field- and current-induced switching. LTspice simulations demonstrated agreement with numerical results, confirming the robustness and fidelity of the model.
@@
This work provides a validated physics-based equivalent circuit model that accurately captures macrospin magnetization dynamics, offering an efficient framework for circuit-level analysis of MTJ devices. The results presented here support the development of next-generation non-volatile memory and advanced computing technologies based on spintronic elements.